The length of a rectangle is 7 less than twice the width. the perimeter is 61. Find the dimensions.
Answer by checkley71(8403)

L=2W-7
P=2L+2W
61=2(2W-7)+2W
61=4W-14+2W
61=6W-14
6W=61+14
6W=75
W=75/6
W=12.5 ANSWER FOR THE WIDTH
L=2*12.5-7
L=25-7
L=18 ANSWER FOR THE LENGTH.
PROOF
61=2*12.5+2*18
61=25+36
61=61